Suppose the bank of namibia undertakes an open market purchase of securities of so that the monetary base increases by $5 billion. By how much will the quantity of money change
1
Expert's answer
2020-08-20T10:17:36-0400

Solution:

Current Monetary base = "120+69=189 \;billion"

Quantity of money in the economy (Money Supplier) ="120+60+69 = 249\;billion"


Money Multiplier = "\frac{Change\;in \;Money\;Supply}{Change\;in\;Monetary \;Base}"


"\frac{249}{189} =1.38"


This means that for every $100 change in monetary base would change the quantity of money by about $138


Suppose the Monetary base is increased by $5 billion, then:

New Monetary base value will be = "\$189 +\$5 = \$194 \;billion"


Money change = "\frac{249}{194}=1.28"


This means that for every $100 change in monetary base would change the quantity of money by about $128

Therefore the quantity of money will change by $128